1.    A normal distribution of BMCC MAT51 scores has a standard deviation of 1.5. Find the z-scores corresponding to each of the following values:

 

a.    A score that is 3 points above the mean.

 

b.    A score that is 1.5 points below the mean.  


1
Expert's answer
2020-08-10T17:14:05-0400


Given : sd = 1.5


formula:


Z= "\frac{x-mean}{sd}"


Solution:

a.

When score is 3 points above the mean , x-mean= 3

so Z= "\frac{3}{1.5}" = 2


b.When score is 1.5 points below the mean , x-mean= -1.5

so Z= "\frac{-1.5}{1.5}" = -1